OpenStudy (anonymous):

PLEASEEEEE HELP I NEED CORRECT ANSWERS!!! :( Aaron is 15 centimeters taller than Peter, and five times Aaron's height exceeds two times Peter's height by 525 centimeters. The system of linear equations that relates Aaron's height (x) and Peter's height (y) is given by (blank) . Aaron's height is (blank) centimeters, and Peter's height is (blank) centimeters

OpenStudy (anonymous):

for the first blank the options are..... x-y = 525 and 5x-2y=15 x-y=-15 and 5x -2y=510 y-x=15 and 5x-2y=525 x-y=15 and 5x-2y=525

OpenStudy (anonymous):

for the second blank the options are... 145 150 165 180

OpenStudy (anonymous):

for the third blank the options are.... 145 150 165 180

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@acxbox22

OpenStudy (anonymous):

@DDCamp

OpenStudy (anonymous):

@rishavraj

OpenStudy (anonymous):

@Loser66

rishavraj (rishavraj):

let Aaron's height be = x peter's height be = y soo x - y = 15

rishavraj (rishavraj):

five times Aaron's height = 5x twice peter's height = 2y so 5x - 2y = 525

OpenStudy (anonymous):

so that would be the first blank right?

OpenStudy (anonymous):

@rishavraj how would i figure out their heights

rishavraj (rishavraj):

just solve the two equations to get "x" and "y" ....

rishavraj (rishavraj):

@madeintokyo

OpenStudy (anonymous):

so aarons height is 165 and peters height is 150? @rishavraj